# Fixing Code Generation Errors for Large Language Models

### v1

@@ VII Conclusion
In this study, we conducted ten rounds of tests on 14 LLMs (such as codeLlama-Python, WizardCoder-Python, and GPT-3.5-turbo) using the HumanEval dataset. Through manual analysis of the test results, we found that these LLMs achieved an average of 84.07% of their reported performance, demonstrating high reproducibility. We also investigated the relationship between Pass@1 results, model inference time, and model parameter size. Our analysis revealed a positive correlation between Pass@1 results and model parameter size, while no significant correlation was observed between inference time and parameter size. Subsequently, we performed an in-depth analysis of errors in the test results, extracting and categorizing 12,837 errors into 14 types. Through in-depth manual analysis, we ultimately identified 19 specific causes leading to these errors. Based on established evaluation criteria, we assessed whether each error cause could be automatically fixed, leading to our development of LlmFix, a method designed to address code generation errors. Evaluation of LlmFix on both HumanEval and MBPP datasets showed that it effectively resolved almost all automatically fixable errors. LlmFix can improve the performance of 14 LLMs by an average of 9.5% and 5.4% on HumanEval and MBPP, respectively.

### v2

@@ 7. Conclusion
In this study, we conducted ten rounds of testing on 14 open-source and closed-source foundation LLMs (e.g., CodeLlama-Python and GPT-3.5-turbo) using the HumanEval dataset to evaluate their reproducibility. Through a detailed manual analysis of the test results, we found that these LLMs achieved an average of 84.1% of their reported performance, indicating a high level of reproducibility. Subsequently, we performed an in-depth analysis of the test results, obtaining and categorizing 12,837 errors into 14 distinct types. After a thorough manual investigation, we ultimately identified 19 specific root causes behind these errors. Based on established evaluation criteria, we assessed whether errors caused by each root cause could be directly fixed, leading to the development of LlmFix—a rule-based static analysis method designed to fix code generation errors. Evaluations of LlmFix were conducted from two perspectives: its performance on error fixing tasks and its impact on improving function-level code generation tasks. For error fixing performance, we constructed a new dataset, LlmErrorEval, based on the 12k+ collected errors to evaluate the performance of fixing methods. Experimental results show that LlmFix achieves a fix rate of 17.1%, outperforming the state-of-the-art LDB by 8.9%. Integrating LlmFix with LDB boosts the fix rate to 24.9% with minimal extra computation overhead. For code generation improvements, evaluations of LlmFix on both the HumanEval and MBPP datasets demonstrated its effectiveness, successfully resolving all directly fixable errors and improving the Pass@1 performance of 14 LLMs by an average of 7.5%.